OpenStudy (anonymous):

Your piggy bank has a total of 18 coins in it; some are dimes and some are nickels. If you have a total of $1.40, how many nickels do you have? 14 12 10 8

OpenStudy (king):

let dimes be x , nickels are y, now x+y=18 also, value(x)+value(y)=1.4$ 10x+5y=140 2x+y=28 now subtract, x=10 y=8

OpenStudy (king):

10 dimes 8 nickels

OpenStudy (anonymous):

Thank's!